# Multiscale Causal Learning

### v2

@@ Abstract
Biological intelligence is more sample-efficient than artificial intelligence (AI), learning from fewer examples. Here we answer why. Given data, there can be many policies which seem “correct” because they perfectly fit the data. However, only one correct policy could have actually caused the data. Sample-efficiency requires a means of discerning which. Previous work showed sample efficiency is maximised by weak-policy-optimisation (WPO); preferring policies that more weakly constrain what is considered to be correct, given finite resources. Biology’s sample-efficiency demonstrates it is better at WPO. To understand how, we formalise the “multiscale-competency-architecture” (MCA) observed in biological systems, as a sequence of nested “agentic-abstraction-layers”. We show that WPO at low levels enables synthesis of weaker policies at high. We call this “multiscale-causal-learning”, and argue this is how we might construct more scale-able, sample-efficient and reliable AI. Furthermore, a sufficiently weak policy at low levels is a precondition of collective policy at higher levels. The higher level “identity” of the collective is lost if lower levels use an insufficiently weak policy (e.g. cells may become isolated from the collective informational structure and revert to primitive behaviour). This has implications for biology, machine learning, AI-safety, and philosophy.
keywords:
multiscale competency, AI, causal learning, sample efficiency
@@ 1 Introduction
Intelligence is often understood as predicting the outputs of a system, given inputs [1, 2, 3]. In that frame, learning is the act of leveraging data [^1] A set of inputs and outputs known to correspond to a particular system. We also refer to this as training or historical data. to improve predictive accuracy. It is an optimisation problem. The cause of data is the policy [^2] By policy, we just mean a constraint on outputs and inputs. In other words, a policy is a declarative definition of behaviour. There is an equivalence between imperative and declarative programs [4] so intuitively, consider a policy to be a program that dictates behaviour. We use the word policy for the sake of consistency with related work and to avoid having to use two different terms for the same thing. We do not mean only the more narrow definition of policy used in reinforcement learning. which originally generated it. To borrow a philosophical term [5], it is the policy which was the original “causal intermediary” between outputs and inputs. In machine learning, training data are sets of examples of goal directed behaviour; an ostensive definition of how to complete a task. A learning algorithm synthesises a policy which “fits” training data. Intuitively, any policy that fits is “correct” for that training data. However, the fact that a policy is correct with respect to data is not enough information to determine whether that policy was the cause of that data. Data could be the result of any number of policies, and if one wishes to accurately predict outputs given inputs that aren’t in the training data, then one must identify that policy which originally generated the data. The true cause. The closer one’s policy is to the cause, the more accurately one will be able to predict outputs given new inputs one hasn’t seen before. Conversely, the better one is at identifying the likely causes of data, the less data one will require to learn how to complete the task. The less data one requires, the more “sample efficiently” one learns.
1.1 Active Inference is Not Enough
Active inference [6, 7] formalises how organisms learn and adapt. It describes living systems in terms of a machine learning approach called variational Bayes. Like other machine learning approaches, this treats learning as a process which synthesises a policy which happens fit data. A policy that is correct relative to historical data, but may not be the true cause of that data. Yet biological intelligence (BI) is far more sample efficient than artificial intelligence (AI). Organoids learn far quicker than deep reinforcement learning agents in silico [8]. Yes, a robust AI will eventually learn cause and effect [9], but for now biology remains for more efficient. How is it that BI is so much more sample efficient than AI? This suggests there is more to the picture than active inference. BI is not just synthesising a policy which “fits” historical data; it is going a step further to figure out which among correct policies are likely to have caused historical data. How?
1.2 Weak-Policy-Optimisation is Causal Learning
Previous work already demonstrates how optimal sample efficiency [10] and causal learning [11] are achievable. Every policy implies a constraint on possible outputs and inputs; a constraint on how many possible tasks [^3] A task means a set of inputs and a corresponding set of outputs. A task is an ostensive definition of a goal or purpose. This is formally defined later in the paper. will be completed by this policy. A more specific policy completes fewer possible tasks, while a policy that implies a weaker constraint completes more possible tasks [12]. Past research into causal learning and sample efficiency found that the correct policies most likely to have caused historical data imply the weakest [^4] Formal definitions for policy weakness are given in the next section. constraint [10, 11] [^5] See Bennett’s Razor and related proofs in [10] for sample efficiency, and [11] for causality arguments.. Experiments demonstrated that a preference for weak policies correctly identifies the cause of historical data at $110-500\%$ the rate of a complexity based induction system [10]. Gradient descent does not optimise for weak policies, which is one reason why deep learning is so sample inefficient [13]. These results are further supported by the contravariance principle proposed by Cao and Yamins [14]. They argue tasks that more tightly constrain the space of correct policies [^6] A larger, more varied set of inputs and outputs implies a smaller set of correct policies. make it easier to identify the policy that generated the inputs and outputs. This supports the claim that weaker policies are more likely to be the cause of behaviour, because a weaker policy is more likely to be a correct policy for tasks that more tightly constrain the set of correct policies. Given the sample efficiency exhibited by organoids [8], it would appear biology is far better at optimising for weak policies than current approaches to machine learning. Causal learning via weak-policy-optimisation [^7] Meaning learning using the weakness proxy as formally defined in the next section and in [3] (WPO) is clearly possible in silico, but existing implementations [10] are computationally expensive. One goal of our research is to better understand how the efficiency we observe in biology might be engineered.
1.2.1 How Weak Policy Optimisation Constructs Causal Identities
Previous work also established that causal learning requires one construct rather than presuppose variables and objects [11]. Causal identities are learned by constructing $X$ such that $X$ causes $Y$, not by presupposing the environment is divided into objects including $X$ and then intervening to test if $X$ causes $Y$ [15]. By optimising for policies that express “weak” constraints on behaviour, otherwise meaningless primitives888e.g. Declarative programs the truth value of which can be known, but for the truth conditions can never be known because environmental states are observable only partially and indirectly [3, 16, 15]. or very low level abstractions may be clustered together to form “causal identities” for higher level abstractions, like variables [11]. The “do” operator which is used to distinguish intervention from observation in causal inference problems can be substituted by introducing variables for interventions [17]. That is why one can learn cause and effect via WPO [11], by constructing causal identities for interventions.
1.2.2 How Causal Identities Relate to Abstraction Layers
The construction of causal identities is a process by which an environment can be broken up into abstract classifications that are inherently meaningful (motivated and relevant to a goal). In philosophical terms, if one is motivated by attraction and repulsion from “contentless” [18] states [^9] Contentless meaning categorised, value-less, monolithic states with no discernible internal structure., then WPO is a means by which those states can be reduced to objects or events which cause the affect which motivates an organism [15]. In computational terms, this amounts to constructing an abstraction layer that is most fit for a purpose [3, 16]. For example, the x86 instruction set architecture (ISA) has hard coded interpretations of certain sorts, such as numbers and arithmetic operations [^10] These amount to policies by the formal definition we use [3, 16]. As a result, the ISA can more efficiently address some problems than others. It is for this reason field programmable gate arrays exist as an alternative to static ISAs like x86; so that specific architectures can be constructed to more efficiently address specific problems.
1.2.3 Static Abstraction Layers and Computational Dualism
The ability to construct causal identities is easily hindered by the presupposition of static abstractions [11, 3]. This is because one cannot construct weaker abstractions if one is forced to use a static set of high level abstractions, like numbers and arithmetic operations. This is symptomatic of problem called “computational dualism[3]. Where 16th century Cartesian interactionist dualism proposed separate mental and physical substances interacting through the pineal gland, computational dualism is the concept of software distinct from hardware. For example, Hinton described software as “immortal”, existing infinitely many times across devices [19]. According to the argument against computational dualism [3], immortal computations do not exist. There are only finitely many instances of mortal computations. Intuitively, when one loads software onto a computer, no object moves into the computer. All that happens is that the state of hardware changes. Addressing computational dualism revealed weakness to be the confounding variable that correlates simplicity with generalisation [16, Prop. 2], allowing objective claims to be made regarding theoretical intelligent agents [3], addressing the subjectivity problems of complexity based induction systems in interactive settings [20]. This is all relevant to us because static abstraction layers are based on computational dualism. They seem to suggest software is something other than the state of hardware. Something interchangeable. Move-able. This means static abstraction layers are necessarily one size fits all, rather than optimised for particular purpose. The inefficiency introduced at the abstraction layer affects everything built within it, including any software based machine learning systems. Intuitively, an abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias. Presupposing the world is formed of numbers, objects and operations of a specific sort limits which cause and effect relations can efficiently be learned. The sample efficiency exhibited by biology suggests it uses a more versatile alternative to static abstraction layers based on computation dualism; one more suitable for WPO.
1.3 Multiscale Competency Architecture
@@
Biology is slowly becoming a matter of engineering. It is now possible to create worms with two heads [21], tadpoles with eyes on their backs that see [22], and biological “xenobots” that perform simple tasks [23]. As biology becomes more tractable, the similarities between biology and human artifice (e.g. computers) have become more obvious. Cells are now understood to be an agentic material [24] for the purpose of engineering. They are but one scale in the “multiscale competency architecture” (MCA) of biology. In an MCA lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [25]. Adaptation takes place at every scale. For example, cells have a “perceptual field”, and can co-operate with other cells in their perceptual field. This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [26]. Cancer happens when this fails [27]. Cells isolated from the informational structure of the collective lose their high level identity, and pursue lower level goals [28]. Biological collectives exhibit high level goal directed behaviours far beyond the capability of any individual part [29]. Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem. In other words, there is a great deal of evidence [26] to suggest that human intelligence is nothing more than the high level goal directed behaviour of a collective of cells. If we want sample-efficient AI, then we should mimic an MCA.
2 Materials and Methods
To summarise what has been covered so far:
  1. 1.
    Our goal is to understand why BI is so much more sample efficient (able not just to synthesise a policy that “fits” data, but to infer which among such correct seeming policies originally generated or caused the data).
  2. 2.
    As sample efficiency is maximised by WPO, our purpose is to understand how biology is able to synthesise weak policies so effectively.
  3. 3.
    A promising avenue appears to be the MCA [25] used by biological systems, instead of the static abstraction layers used in conventional computers (which exhibit the shortcomings of computational dualism [3]).
Hypothesis: The agentic, nested structure of MCAs allows more effective and efficient WPO than would be possible within a static layer like x86.
2.1 Agentic Abstraction Layers
@@
To begin to understand how this might be possible, we must formalise MCAs in rigorous, mathematical terms. High level goal directed behaviour can emerge from distributed systems like swarms of cells, but it is extremely difficult to precisely control this sort of emergent high level goal directed behaviour. Work by Ciaunica et. al. stressed the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [30]. Taking our cue from that approach, we propose formalising MCAs as nested and agentic abstraction layers. These abstraction layers are descriptions of behaviour, so we need not concern ourselves with the exact details of the system. It could be like biology, or something else. Our primary concern is how we might engineer and control efficient causal learning. A non-agentic, static abstraction layer passively interprets the abstraction layers above; like how a CPU based on the x86 instruction set architecture (ISA) interprets the behaviour of a python interpreter. By an agentic abstraction layer, we mean one that has its own agenda. It adapts. It compels the layers above to serve its own ends. Higher levels are compelled by lower to learn a policy “aligned” with theirs. Lower levels change their interpretation to enact [31] policy through higher levels, and are constrained by the resulting higher level behaviour. Information passes in both directions. Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate the pursuit of higher level goals.
2.2 Definitions
We begin with a formalism [3] of enactive cognition [31]. We will use it to formalise MCAs in the next section. More extensive explanations of these definitions are given elsewhere [16, 3], and we have endeavoured to make key aspects of this paper understandable without the formalism; readers who wish to should simply skip to the next section.
Definition 1. [environment].
@@ 2.2 Definitions
  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all declarative programs (the powerset of $\Phi$).
@@
  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t. $\phi\in\bigcap l$. By an aspect of the environment we mean an aspect $l$ of any state, s.t. $\bigcap l\neq\emptyset$. We say an aspect of the environment is realised by state $\phi$ if it is an aspect of $\phi$.
Definition 2. [abstraction layer] .
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite unless stated otherwise [^11] This is because a bounded system can contain only a finite amount of information [32] in a spatially extended environment like our own [16].. If $\mathfrak{v}=P$, then we say that there is no abstraction.
@@
(notation) $E$ with a subscript is the extension of the subscript121212e.g. $E_{l}$ is the extension of $l$..
(intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer. The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table.
Definition 3. [$\mathfrak{v}$-task] For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
@@
(intuitive summary) To reiterate and summarise the above:
@@
  • An output is a completion of an input [def. 2.2].
@@
2.2.1 Learning and inference definitions:

### v3

# Are Biological Systems More Intelligent Than Artificial Intelligence?

@@ Abstract
Is a biological self-organising system more ‘intelligent’ than an artificial intelligence? If so, why? We frame intelligence as adaptability, and explore this question using a mathematical formalism of enactive causal learning. We extend it to formalise the multilayer, multiscale, bottom-up distributed computational architecture of biological self-organisation. We then show that this architecture allows for more efficient adaptation than the static top-down interpreters typically used in computers. To put it provocatively, biology is more intelligent because cells adapt to provide a helpful inductive bias, and static interpreters do not. We call this multilayer-causal-learning. However it inherits a flaw of biological self-organisation. Cells become cancerous when isolated from the collective informational structure, reverting to primitive transcriptional behaviour. We show that, in the context of our formalism, failure states like cancer occur when systems are too tightly constrained by the abstraction layer in which they exist. This suggests control should be distributed (bottom-up rather than top-down) to ensure graceful degradation. We speculate about what this implies for systems in general, from machine learning hardware to human organisational and economic systems. Our result shows how we can design more robust systems and, though theoretical in nature, it lays a foundation for future empirical research.
@@ 1 Introduction
Biological systems self-organise, meaning their parts and the environment interact to produce spatiotemporal order [1, 2, 3, 4, 5, 6, 7]. They interact with their environment to achieve homeostatic and reproductive goals. Some proactively learn [8], storing information about the past in order to anticipate future needs. To make accurate predictions in environments of sufficient complexity, a system must be able to learn cause and effect [9, 10, 11]. For our purposes we consider the ‘intelligence’ of a system to be a measure of the resources it needs to correctly identify cause and effect [12, 13, 14, 15, 16] [^1] There are many definitions of intelligence, the merits of which are beyond the scope of this paper. Ours is primarily informed by [12], which is adapted from Legg and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’. The more efficient the learning system, the wider the range of environments in which it can satisfy goals. It is not possession of skills, but an ability to acquire skills with limited resources. Being an embodied formalism, this also accounts of energy and other resources consumed.. It is a measure of how efficiently one adapts [^2] Primarily sample and energy efficiency. A ’sample’ is information used to learn. When a teacher provides an example of a math problem being solved, they are providing a sample. If student A requires 4 samples to understand a concept, and student B requires 5 samples, then A is more sample efficient than B..
Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) of those that learn [17, 18, 19, 20]. Contemporary approaches to machine learning are notoriously inefficient [21]. They need vast quantities of data and energy learn tasks most humans can pick up after one or two examples. Even organoids, small collectives of cells grown in a lab, appear to be more ’intelligent’ in this sense than contemporary machine learning [22]. Why is this the case? Are biological systems more intelligent, or is something else going on?
Some suggest that the relative inefficiency of ML may be down to the fact that biological systems are active inference systems, and machine learning systems are typically not [23, 6, 24]. Active inference is much more efficient [25, 26, 27], but it may only be part of the picture. As of 2024 active inference systems [^3] In particular the models submitted by Jack Cole and his employer, Minds AI. applied in the context of the ARC benchmark [14] [^4] For additional information, see https://arcprize.org for sample efficient adaptation still under-perform humans by a substantial margin [^5] Though it should also be noted that the Minds AI active inference system is one of the leading ARC submission for 2024, excluding OpenAI’s o3 which may or may not be an active inference system (the details have not been released at this time), and which is disqualified from ARC because it does not meet the computational efficiency requirements.. Others hold that the reason biological systems appear to be more intelligent is because of an ‘inductive bias’ [14]. This means the system embodies hard-coded foreknowledge of the environment, which predisposes it to efficient adaptation. If an inductive bias were the only reason for the discrepancy between biological self-organisation and machine learning, that would suggest biological systems are not more intelligent than artificial intelligence, but more knowledgeable.
To reiterate, we seek to explain why biological self organisation seems to be more sample (and energy) efficient than contemporary machine learning. Is there something biological systems do that allows them to be more ’intelligent’ than artificial intelligence in its contemporary form? We suggest the answer lies in how biological self organisation distributes control at a different scales, and levels of abstraction. To do so we must first give some background on causal learning which informs our hypothesis.
1.1 Causality
Cause and effect are typically represented as an acyclic graph.
causeeffect
Complex environments often complicate the ability to identify causes. For example, confounded events may seem to cause one another when in fact they are merely correlated. This can lead one to draw erroneous conclusions about cause and effect.
not the causeeffectconfounder
By actively intervening in the environment a system can experiment to isolate causes. To do so, a system must be able to differentiate between events it passively observes and events it has caused [10, 9, 28]. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relation between them. This works well in theory for artificial intelligence [29], but requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges between nodes in the graph), and then learn the variables (the abstract concepts). This needs only attraction and repulsion from physical states [30], otherwise known as valence. We take this latter approach.
We can frame causal learning in computational terms. Learning nodes in a causal graph is like figuring out which computer program was used to generate a sequence. Assume we have fed a set of inputs into a computer, and for each input the computer responded with a particular output. The computer is running a ’policy’ program which maps inputs to outputs [^6] We’ll call all such programs policies for the sake of this paper. This should not be confused with the more specific notion of a policy from reinforcement learning [19].. That policy is the cause of the input-output history we’ve observed so far. We want to know the policy that caused historical outputs so that we can use it to predict future outputs. The problem is that more than one policy exists which could have produced the history we’ve observed, and each of those policies produce different outputs when we give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have been the true underlying cause. If we wish to know which policy actually caused the outputs, we need way to narrow down the possibilities even further.
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [31]. One way is to can compare their form, and choose programs which are simpler. However, in an interactive setting simplicity is an unreliable indicator of plausibility [32, 31]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which we do not have historical data. Adaptive systems frequently satisfy the same functional constraints by different means [33], so we can compare different sorts by the constraints they satisfy. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the greater the variety of future behaviours it permits. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker of the two [16]. In theoretical terms the policy most likely to have caused behaviour is the weakest, and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [34]. This has been used to establish local upper bounds on embodied ‘intelligence’ [12].
ML systems do not typically optimise for either weaker constraints on function, or simplicity of form. This at least partially explains their relative sample inefficiency [35, 34]. However, even when ML systems are modified to optimise for weakness, their capacity to do so limited by the “abstraction layer”. An abstraction layer is like a causal isolator [36] between a software “mind” and its environment. For example, the Python programming language is interpreted by the Python interpreter. In the end, Python gets translated into machine code, which hardware like an Apple M2 system on a chip then interprets to physical effect based on the ARM instruction set architecture. An abstraction layer is just physical embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, and thus different upper bounds on embodied ’intelligence’ [12]. In other words, physical embodiment imposes severe limitations, and the behaviour of software is determined entirely by the hardware and software abstraction layers in which it runs. Intelligence is primarily a matter hardware, not software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [37, 38, 39].
Put another way an embodied policy is a physical structure, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. A more specific tool completes fewer possible tasks, while a tool that implies a weaker constraint completes more possible tasks [21]. By preferring weaker policies, one avoids unnecessary assumptions and so adapts as efficiently as possible, meaning the weakest policies complete a ‘wider variety of tasks’ than all others [^7] As an informal epistemological razor, this means “explanations should be no more specific than necessary”. Intuitively, this is a similar idea to the principle of maximum entropy [40, 41], but it accounts for the subjectivity of complexity and the teleological nature of language [31]. For further intuition, a corollary to the idea that a weaker policy is more likely to be the cause of data is the inventors paradox, or equivalently the contravariance principle proposed by Cao and Yamins [42]. They argue tasks that more tightly constrain the space of correct policies make better training sets. A larger, more varied set of inputs and outputs implies a smaller set of correct policies. That makes it easier to identify the policy that generated the inputs and outputs. This supports the claim that weaker policies are more likely to be the cause of behaviour, because a weaker policy is more likely to be a correct policy for tasks that more tightly constrain the set of correct policies. See proofs 1 and 2 for more details.. To complete a ‘wider variety of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses.
That biological self-organising systems adapt efficiently suggests they are able to embody weaker policies than today’s ML systems. They attain higher local upper bounds on embodied intelligence [12]. This suggests they have a means of optimising not just the choice of policy, but the embodied abstraction layer in which that policy exists.
In a typical computer, the abstraction layers are static. For example, a Python interpreter does not learn and adapt to suit the program it is running. An abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias. When we say that biological systems can optimise the abstraction layer, this is like saying they can “move” this window to learn a different inductive bias. The question is, what is the abstraction layer in the context of a biological self-organising system?
@@ 1.1 Causality
We suggest the biological equivalent of abstraction layers are the layers of a “multiscale competency architecture” (MCA). In an MCA lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [43]. Adaptation takes place at every scale, and cells are like an agentic material [44]. Cells have a “perceptual field”, and can co-operate with other cells in their perceptual field. This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [45]. Each cell pursues goals, and then ever larger groups of cells pursue collective goals at ever larger scales. Biological collectives exhibit high level goal directed behaviours beyond the capability of the individual parts [46]. Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem. In computational terms this is a distributed system, like a von Neumann architecture [33].
Our somewhat provocative claim is that biological self-organising systems are inherently more intelligent than artificial intelligence because their distributed, multilayer, multiscale architecture allows for higher local upper bounds on ’intelligence’. This is because it can shift abstraction layers to convey helpful inductive biases. Our results, though purely theoretical, lay a foundation for future empirical research.
In section 2 introduce the formalism. In section 3 then argue this formalism represents an MCA and explain how. In the context of that formalism, we will formally show that an MCA allows for higher local upper bounds on embodied ‘intelligence’. We call this multilayer-causal-learning (MCL). In section 4 we explore how MCL may inherit the flaws of biological self-organising systems as well as their increased adaptability. We formalise a failure state analogous to cancer, and show that this state is more likely when a system is more tightly constrained. Finally, we relate these results to systems in general. We speculate about the possible implications, from more efficient hardware for machine learning to improved human organisational systems.
2 Formalism and Proofs
@@
As part of our argument we attempt to describe MCAs in mathematically precise terms. Recent research has stressed the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [47]. Taking our cue from that approach, we will formalise MCAs as a series of nested, adaptive, agentic abstraction layers. Enactivism frames cognition as an interaction between a self-organising system and its environment environment [48, 49, 50]. We merely extend this process inwards, to formalise the interactions within the system itself. A non-agentic, static abstraction layer passively interprets the abstraction layers above. By an agentic abstraction layer, we mean one that has its own agenda. It adapts and compels the layers above to serve its own ends, while they each serve their own increasingly specific goals within the bounds of what lower levels permit. Higher levels are compelled by lower to learn a policy “aligned” with lower levels. Lower levels change their interpretation to enact [48] policy through higher levels, and are constrained by the resulting higher level behaviour. Information passes in both directions. Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate the pursuit of higher level goals.
Except for the last definition, all the of the definitions that follow are from preceding work [12, 31]. They formalise self-organisation in terms of constraints, like boundary conditions, as this lends itself to enactive depictions of intelligence [51]. We have endeavoured to make key aspects of this paper understandable without the formalism. Readers who wish to should simply skip to the next section.
{adjustwidth}
11.6pt11.6pt
Definition 1 (environment).
@@
  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all programs (the powerset of $\Phi$).
@@
  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t. $\phi\in\bigcap l$. By an aspect of the environment we mean an aspect $l$ of any state, s.t. $\bigcap l\neq\emptyset$. We say an aspect of the environment is realised [^8] Realised meaning it is made real, or brought into existence. by state $\phi$ if it is an aspect of $\phi$.
Definition 2 (abstraction layer).
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite.
@@
(notation) $E$ with a subscript is the extension of the subscript999e.g. $E_{l}$ is the extension of $l$..
$L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer. The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table. Intuitively a sensorimotor system is an abstraction layer. Likewise, a computer (each statement asserting a state of the computer, or a part thereof).
{adjustwidth}
11.6pt11.6pt
Definition 3 ($\mathfrak{v}$-task).
For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
@@
To reiterate and summarise the above:
@@
  • An output is a completion of an input [def. 2].
@@
2.1 Learning and inference definitions